# Fair-Gate: Fairness-Aware Interpretable Risk Gating for Sex-Fair Voice Biometrics

### 11 Mar 2026 v1

@@ I Introduction
Inspired by causal fairness formulations [16], we distinguish (i) the causal effect of sex on speech acoustics (e.g., $F_{0}$ and formants) from (ii) dataset-induced correlations between sex and speaker identity in the training data. As illustrated in Fig. 1, standard discriminative training can exploit such correlations as demographic shortcuts, shifting score distributions differently across groups and producing unequal subgroup error rates under a shared threshold. In this work, fairness is evaluated across binary proxy sex groups inferred from speech by a frozen classifier. Accordingly, we use the term sex to refer to these proxy labels, and we do not claim to measure self-identified gender.
@@ II-A Encoder
Given an input log-Mel spectrogram $x\in\mathbb{R}^{F\times T}$, the encoder $E(\cdot)$ produces frame-level features in $\mathbb{R}^{C\times T}$. For mini-batches of size $B$, we denote the resulting tensor by $\mathbf{U}\in\mathbb{R}^{B\times C\times T}$.
@@ II-B1 Mask computation and routing
Both branches retain dimensionality and satisfy $\mathbf{U}_{\mathrm{id}}+\mathbf{U}_{\mathrm{sex}}=\mathbf{U}$ element-wise. This makes routing additive and lossless at the feature level, allowing Fair-Gate to learn where information should be represented rather than forcing identity- and sex-related cues into fixed disjoint subspaces.
@@ II-B2 Gate regularization
To avoid degenerate routing (e.g., collapsing almost all features to one branch or producing ambiguous allocations), we regularize the mask $A$ with two complementary terms: a routing-mass control term $\mathcal{L}_{\mathrm{cap}}$, which prevents branch collapse, and a saturation term $\mathcal{L}_{\mathrm{sat}}$, which encourages confident near-binary routing decisions.
@@ II-C5 Risk Variance Equalization
Risk Extrapolation (REx) acts on speaker-classification risk rather than on sex predictability. In our setting, the proxy sex groups define the environments over which risk variation is measured. Let $\mathcal{E}=\{\mathrm{M},\mathrm{F}\}$ and let $\mathcal{D}_{e}$ denote the subset of training samples assigned to proxy group $e$. The per-group speaker-classification risk is
$\mathcal{R}_{e}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{e}}\mathrm{CE}\big(h_{\mathrm{spk}}(z_{\mathrm{id}}(x)),y\big),\qquad\bar{\mathcal{R}}=\tfrac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathcal{R}_{e}.$ (8)
@@
Intuitively, if the speaker classifier relies on group-specific shortcuts, speaker-classification risk will differ systematically across proxy sex groups. Penalizing the variance of $\{\mathcal{R}_{e}\}$, therefore, encourages the model to rely on speaker evidence that transfers more uniformly across groups.
@@ II-D Overall Training Objective
TABLE II: Results on VoxCeleb1 (O/E/H). We report EER [%], minDCF, and sex disparity measured by GARBE at the threshold $\tau_{1\%}$.
@@
TABLE III: Expanded ablation results on Vox1-H. We report EER [%], $\mathrm{GARBE}(\tau_{1\%})$, and subgroup-specific FNMR/FMR [%] at the same shared threshold $\tau_{1\%}$. Here, $\tau_{1\%}$ is first determined on pooled Vox1-H trials and then reused unchanged for both proxy sex groups. Bold indicates the best value in each column based on the underlying unrounded scores.
@@
Main, Full (r=0.005r{=}0.005) 2.25 0.07 0.96 1.07 3.80 4.49
Main, Full (r=0.02r{=}0.02) 2.32 0.07 0.95 1.08 3.98 4.71
@@ III-B Metrics and baselines
    False match rate (FMR): proportion of zero-effort impostor attempt samples falsely declared to match the compared non-self template;
    False non-match rate (FNMR): proportion of genuine attempt samples falsely declared not to match the template of the same characteristic from the same user supplying the sample;
@@
$\begin{gathered}\mathrm{DCF}(\tau)=\\ C_{\mathrm{FNMR}}\,P_{\mathrm{target}}\,\mathrm{FNMR}(\tau)+C_{\mathrm{FMR}}\,(1-P_{\mathrm{target}})\,\mathrm{FMR}(\tau),\\ \\ \mathrm{minDCF}=\min_{\tau}(\mathrm{DCF}(\tau)),\end{gathered}$ (11)
@@
While identical thresholds are used for male and female score distributions for all experiments, different thresholds $\tau$ are computed according to the system evaluated (baselines and proposed) and the metric. For EER assessment, $\tau$ corresponds to the decision threshold where the system achieves $FMR=FNMR$. For the computation of $minDCF$, $\tau$ is selected such that $DCF$ is minimised. Finally, for GARBE, the threshold is fixed at $\tau=$, such that $FMR(\tau)=10^{-2}$, which we indicate in the following with $\mathrm{GARBE}(\tau_{1\%})$.
@@
All methods are trained on the VoxCeleb2 development set and evaluated on VoxCeleb1 using the original (O), extended (E), and hard (H) trial protocols. To ensure a fair comparison, we follow identical training data splits, preprocessing steps, and evaluation metrics across all systems.
@@ IV-B Ablation study on VoxCeleb1-H
The remaining ablations indicate that the other components play more limited or supporting roles. Removing REx still degrades the trade-off, increasing EER from 2.25% to 2.55% and GARBE from 0.07 to 0.08, while also raising subgroup FMRs from 3.80% / 4.49% to 4.54% / 5.60%, which shows that risk equalization contributes meaningfully under the shared operating point. By contrast, removing the adversarial term changes fairness only marginally but slightly worsens utility (2.25% $\rightarrow$ 2.27%), suggesting that adversarial invariance is not the primary source of subgroup-gap reduction. Finally, removing $L_{\mathrm{sat}}$ causes only a mild degradation, and increasing the REx weight from $r=0.005$ to $r=0.02$ does not further improve the result, indicating that moderate risk equalization is preferable to stronger regularization in the present setup.

### 01 Apr 2026 v3

@@ I Introduction
Inspired by causal fairness formulations [16], we distinguish (i) the causal effect of sex on speech acoustics (e.g., $F_{0}$ and formants) from (ii) dataset-induced correlations between sex and speaker identity in the training data. As illustrated in Fig. 1, standard discriminative training can exploit such correlations as demographic shortcuts, shifting score distributions differently across groups, and producing unequal subgroup error rates under a shared threshold.
@@
Despite sex and gender terms being commonly used interchangeably, there is a generally, though not universally accepted distinction. We adopt the terminology in [21] and expressly avoid references to gender which refers to socially constructed roles and behaviour. Instead, we refer only to sex, which refers to biological attributes [24].
@@ II-A Encoder
Given an input log-Mel spectrogram $x\in\mathbb{R}^{F\times T}$, where $F$ is the number of Mel-frequency bins and $T$ the number of frames, the encoder $E(\cdot)$ produces frame-level features in $\mathbf{U}\in\mathbb{R}^{C\times T}$, where $C$ denotes the number of feature channels.
@@ II-B1 Mask computation and routing
Both branches retain dimensionality and satisfy $\mathbf{U}_{\mathrm{id}}+\mathbf{U}_{\mathrm{sex}}=\mathbf{U}$ element-wise. This additive decomposition ensures that the gate reallocates information between branches, allowing Fair-Gate to learn where information should be represented rather than forcing identity- and sex-related cues into fixed disjoint subspaces.
@@ II-B2 Gate regularization
To avoid degenerate routing (e.g., collapsing all features to one branch or producing ambiguous allocations), we regularize the mask $A$ through two complementary terms defined below: a routing-mass control term $\mathcal{L}_{\mathrm{cap}}$ and a saturation term $\mathcal{L}_{\mathrm{sat}}$.
@@ II-C5 Risk Variance Equalization
Risk Extrapolation (REx) reduces discrepancies in speaker-classification risk loss across proxy sex groups. Let $\mathcal{E}=\{\mathrm{M},\mathrm{F}\}$. The per-group risk is
$\mathcal{R}_{e}=\mathbb{E}_{(x,y)\sim\mathcal{E}}\mathrm{CE}\big(h_{\mathrm{spk}}(z_{\mathrm{id}}(x)),y\big),\qquad\bar{\mathcal{R}}=\tfrac{1}{|\mathcal{E}|}\sum_{e\in\mathcal{E}}\mathcal{R}_{e}.$ (8)
@@
Intuitively, if the speaker classifier relies on group-specific shortcuts, speaker-classification risk will differ systematically across proxy sex groups. Penalising the variance of $\{\mathcal{R}_{e}\}$, therefore, encourages the model to rely on speaker evidence that transfers more uniformly across groups.
@@ II-D Overall Training Objective
TABLE II: Results on VoxCeleb1 (O/E/H). We report EER [%], minDCF, and sex fairness measured by GARBE at the threshold $\tau_{1\%}$.
@@
TABLE III: Ablation results on Vox1-H. We report EER [%], $\mathrm{GARBE}(\tau_{1\%})$, and subgroup-specific FNMR/FMR [%] at the same shared threshold $\tau_{1\%}$. Bold indicates the best value in each column.
@@
Main, Full 2.25 0.07 0.96 1.07 3.80 4.49
@@ III-B Metrics and baselines
    False match rate (FMR): proportion of the completed biometric non-mated comparison trials that result in a false match;
    False non-match rate (FNMR): proportion of the completed biometric mated comparison trials that result in a false non-match;
@@
$\mathrm{DCF}(\tau)$ $=C_{\mathrm{FNMR}}\,P_{\mathrm{target}}\,\mathrm{FNMR}(\tau)$ (11)
$\quad+C_{\mathrm{FMR}}\,(1-P_{\mathrm{target}})\,\mathrm{FMR}(\tau),$
$\mathrm{minDCF}$ $=\min_{\tau}\,\mathrm{DCF}(\tau).$
@@
While identical thresholds are used for male and female score distributions for all experiments, different thresholds $\tau$ are computed according to the system evaluated (baselines and proposed) and the metric. For EER assessment, $\tau$ corresponds to the decision threshold where the system achieves $FMR=FNMR$. For the computation of $minDCF$, $\tau$ is selected such that $DCF$ is minimised. Finally, for fairness, the threshold for GARBE is fixed at $\tau=1\%$, such that $FMR(\tau)=10^{-2}$, which we indicate in the following with $\mathrm{GARBE}(\tau_{1\%})$.
@@
All methods are trained on the VoxCeleb2 development set and evaluated on VoxCeleb1 using the original (O), extended (E), and hard (H) trial protocols. To ensure a fair comparison, we follow identical training data splits, preprocessing steps, and evaluation metrics across all systems. Unless otherwise stated, all reported experiments use a batch size of 512.
@@ IV-B Ablation study on VoxCeleb1-H
The remaining ablations indicate that the other components play more limited or supporting roles. Removing REx still degrades the trade-off, increasing EER from 2.25% to 2.55% and GARBE from 0.07 to 0.08, while also raising subgroup FMRs from 3.80% / 4.49% to 4.54% / 5.60%, which shows that risk equalization contributes meaningfully under the shared operating point. By contrast, removing the adversarial term changes fairness only marginally but slightly worsens utility (2.25% $\rightarrow$ 2.27%), suggesting that adversarial invariance is not the primary source of subgroup-gap reduction.
@@ V Conclusions
Acknowledgement
This work was supported by the French Agence Nationale de la Recherche (ANR) via the SpeechPrivacy (ANR-23-CE23-0022) and GOOD-BIAS (ANR-25-CE39-6459) projects.